OpenStudy (anonymous):

At the start of the basketball game, the referee tosses a ball for a jump ball. The equation that models the height of the ball is h(t) = -16t2 + 24t + 5 During what time interval is the height of the ball above 9 feet? (Round your answer to the nearest tenth of a second.)

OpenStudy (rituraj):

9 = -16(t^2) + 24t + 5 x = 1.3s, .2s (Using quadratic formula)

OpenStudy (anonymous):

how do you know what x is equal to

OpenStudy (rituraj):

x = (-b+sqrt ((b^2) - 4ac)/2a, (-b-sqrt ((b^2) - 4ac)/2a for a(x^2) + bx +c = 0